[Brahms-l] Shift report 20050510 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue May 10 2005 - 08:01:22 EDT
Shifter leader was rw
Shifters were 


1. 6 hours beam from 2:00am till now. 
   This store is scheduled to be dumped at 9:00am.

2. run 14538 - 14548.
   No pedestal runs

3. No detector problem is realized.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue May 10 08:01:35 2005

This archive was generated by hypermail 2.1.8 : Tue May 10 2005 - 08:01:47 EDT